OpenStudy (anonymous):

Solve the system of inequality: x + y + z = -­1 2x + 4y + z = 1 x + 2y -­ 3z = ­-3

rebeccaxhawaii (rebeccaxhawaii):

elimination method?

OpenStudy (anonymous):

@rebeccaxhawaii Yes!

OpenStudy (anthonyym):

I think you can type them in as a matrix in the graphing calculator and then hit rref.

OpenStudy (anonymous):

@anthonyym unfortunately I don't own a graphing calculator :/

OpenStudy (anonymous):

Use the first 2 equations to get x and z in terms of y. Then substitute those values in 3rd equation to solve for y. Once you have y, get values of x and z from original substitutions. This is probably fairly hard to follow, but here is the actual equation solving, i just did. z = -1 -x -y ---> z = 1 - (2-3y) -y = 1 - 2 + 3y - y = -1 + 2y z = 1+2(8/3) = 19/3 2x + 4y + (-1-x-y) = 1 2x + 4y -1 -x -y = 1 x + 3y -1 = 1 x + 3y = 2 x = 2 -3y = 2 - 3(8/3) = -6 x + 2y -­ 3z = ­-3 2-3y +2y - 3 (-1+2y) = -3 2 -y + 3 - 2y = -3 5 - 3y = -3 y = 8/3

OpenStudy (anonymous):

Thank god, that was really helpful thank you so much!